OpenStudy (anonymous):

4x - 2y = 8 8y = -16x - 32

OpenStudy (shinalcantara):

4x - 2y = 8 ------equation 1 8y = -16x - 32 ------equation 2 you can do either elimination and substitution method. Let's do substitution. Divide the whole equation 2 by 8

OpenStudy (shinalcantara):

what did you get?

OpenStudy (08surya):

multiply equation with 4 and then add it with equation 2 you will get the value of x

jhonyy9 (jhonyy9):

4x-2y=8 8y=-16x -32 this is the equal 16x +8y =-32 divide both sides by 4 and will get 4x +2y = -8 so your exercise wil be in this case 4x -2y = 8 4x +2y = -8 ------------ add them and will get 8x 0 = 0 8x=0 x=0 4*0 -2y = 8 0 -2y = 8 2y= -8 y = -4 x=0 y=-4 hope this is understandably